In the industrial preparation of ammonia, nitrogen and hydrogen gases are are reacted according to the reaction N2(g) + 3H2(g) 2NH3(g). Suppose 466 kg of N2 are mixed with 39.4 kg of H2

How many moles of N2(g) and of H2(g) will remain when the reaction is complete?
1
Expert's answer
2019-09-18T01:06:00-0400

At first, the limiting reactant should be identified. For the purpose, the number of moles of reactants must be estimated.

Mr(N2) = 28 g/mol

Mr(H2) = 2 g/mol

Next, according to the reaction, the number of moles of the reactants are: n(N2) = 466 × 103 g / 28 g/mol = 16.64 × 103 mol and n(H2) = 39.4 × 103 g / 6 g/mol = 6.57 × 103 mol. As a result, hydrogen gas is a limiting reactant and nitrogen gas is in excess. The reaction occurs when equal amounts of both gasses react. As a result, 16.64 × 103 - 6.57 × 103 = 10.07 × 103 moles of nitrogen will not react and will remain while all amount of hydrogen will be used in the reaction.


Answer: 10.07 × 103 mol of N2(g) and 0 mol of H2(g).